# When Large Language Models Meet UAVs: How Far Are We?

### v1

@@ 5.2.1. Practices and Challenges in UAV-LLM integration
Although LLMs show strong potential in auxiliary tasks, their stability, interpretability, and real-time performance have not yet met the high standards required for industrial applications in core tasks like flight control. Therefore, LLMs are currently more applied in low-risk scenarios such as Task Planning and decision support, while their use in high-risk areas like flight control remains highly restricted. One practitioner in the survey pointed out, “I do not think LLMs are needed for tasks inboard the uav. AI can have a place on board for image processing to add sensors. But AI should never be put in direct control of a uav. It is an unnecessary, lazy, and non-robust approach to vehicle control.” This clearly reflects the cautious attitude of the industrial sector towards the use of LLMs in critical UAV control tasks, with the belief that LLMs are more suitable for supporting tasks rather than directly participating in flight control.
@@
The Need of UAV-LLM integration. Additionally, we asked practitioners about the level of need for LLMs in different UAV tasks. The results show that practitioners have a higher demand for Information Input (score 2.42) and Task Planning (score 2.53). This indicates that LLMs have significant application potential in processing user command translation, perception data, and Task Planning. These tasks require high data processing, reasoning, and planning capabilities, and therefore, there is a strong demand for LLMs’ involvement in these areas. In contrast, the demand for Control Execution (score 2.05) is lower, suggesting that in the areas of flight control and command generation, existing technologies are already well-suited to meet the needs, or the current large language model technology cannot be fully relied upon. This indicates that in real-time flight control systems, traditional control algorithms and hardware systems still dominate.
Finding 10: Regarding task demands, Information Input (score 2.42) and Task Planning (score 2.53) received higher scores, and 82.7% of practitioners believe that Onboard Assistance is the primary way to integrate large language models with UAVs. This indicates that, at present, LLMs in the UAV domain are more focused on supportive tasks rather than directly intervening in the flight control system.
5.2.2. Industry-Academia Comparison.
@@
Practitioner’s Perception of the Academia-Industry Gap. To understand the practical application of academic research in industrial settings, we asked practitioners whether they refer to academic research in their actual work and included an open-ended question asking for the reasons. The results show that 69.2% of practitioners indicated that they do refer to academic research. The main reasons for this are as follows: gaining cutting-edge knowledge, verifying the accuracy of information, accelerating development speed, and learning advanced algorithms and technologies. In particular, in the area of algorithm development, academic research provides developers with new ideas and methods, which have actively contributed to improving work efficiency and development quality. For example, one practitioner mentioned, “I am implementing algorithms for paths’ planning, etc., from the academic publications, I also use them during brainstorming of new ideas and to learn state-of-the-art.
Finding 11: 69.2% of practitioners stated that they would reference academic research results in their actual work to stay updated with the latest technologies, which suggests that academic research plays an important role in driving UAV-LLM practices.
@@
Although most practitioners tend to refer to academic research, a small portion of practitioners do not directly refer to academic outcomes. In this regard, we also asked for their views on the differences between academia and industry in the process of integrating UAVs with large models. According to the survey results, 51.9% of practitioners believe that industry is progressing faster than academia. This suggests that while academia is advancing in theoretical research and exploring new technologies, the actual application and implementation in industry are happening at a faster pace. In contrast, 36.5% of practitioners believe that academia is progressing faster than industry. This reflects that, in basic theoretical research and technological innovation, academia’s breakthroughs provide new ideas and directions for the industry. However, because academic research is not always aligned with industrial demands, the practical application of these new technologies progresses relatively slowly. For example, one practitioner stated, “Industry is faster than academia at some things that lead to higher technology readiness levels. Industry tends not to be good at starting low trl development. In my experience, academia often develops things with little to no industrial application.” This highlights the advantage of industry in applying technology, especially in fields with higher Technology Readiness Levels (TRLs), where industry’s pace of advancement is faster. However, academia often conducts research at low TRL stages, leading to a disconnect between certain academic studies and practical industrial applications.
@@
Additionally, 38.5% of practitioners believe that there are differences between academia and industry in the integration methods of LLMs and UAVs. 67.3% of practitioners think that academia and industry focus on different directions. Some practitioners pointed out: “Research - one topic, but deep, with logical knowledge base and understanding, industry - a lot of topics skimmed, practical usage with ‘black-box-magic”’. These statements illustrate that academia typically focuses on deeply solving specific problems and exploring unknowns and theoretical innovations, while industry is more concerned with practical applications, focusing on how to quickly, efficiently, and reliably apply technology.
Table 3. Practitioners’ perceptions on the reasons for differences
@@
Technical Actual projects require mature and stable technology (75.0%) 84.6%
-Maturity Difficulty in actual project development and integration (55.8%)
@@
Solutions to Bridge the Academia-Industry Gap. Furthermore, we asked practitioners to identify the reasons behind these differences. The responses can be categorized into the five challenges listed in  3. Among them, 84.6% of participants identified “Technical Maturity” as the main challenge. This indicates that, unlike academic research, which often emphasizes theoretical innovation, industry places greater importance on mature and reliable technical solutions. Specifically, practitioners pointed to the need to address issues such as “Actual projects require mature and stable technology” (75.0%), “Difficulty in actual project development and integration” (55.8%), and “Immaturity of lightweight large models” (40.4%). These findings reflect the challenges academic research faces in translating results into real-world applications, as industry often encounters complex conditions during technology integration and project development.
Practitioners also noted that industry faces challenges related to “Performance” (48.1%), “Safety” (42.3%), “Cost” (44.2%), and “Other” (57.7%). Corresponding issues include “High computational resource requirements” (46.2%), “Conflicts between latency and real-time requirements” (34.6%), “High computation/hardware costs” (32.7%), "Model robustness and reliability challenges" (30.8%), and “Commercial return on investment” (28.8%). One practitioner emphasized: “Industry wants practical results fast, especially smaller startups, etc. Academia has smaller groups to work with, less money, but great research skills and expert knowledge (even if not always practical).” This highlights the urgent demand for rapid results in industry, particularly in the process of commercialization, which makes the transfer of academic research outcomes into practical industrial applications more challenging.
Finding 12: Practitioners believe that the differences between academia and industry primarily lie in their focus areas (67.3%) and technical requirements (63.5%). The greatest challenge in addressing these differences is technical maturity (84.6%). Based on this, we can infer that although academic research can drive technological innovation, the development and integration of technology in practical projects face higher complexity and challenges. This necessitates consideration of the maturity, stability, and practical feasibility of the technology. This suggests that more attention should be paid to the alignment between academia and industry in terms of technology transfer and real-world application.
@@ 6. Discussion
This section first explains the significance of this study for researchers and practitioners, then summarizes the three main challenges of UAV–LLM applications, and finally presents the validity threats.
6.1. Implications
@@
For researchers. This study reveals the significant differences between academia and industry in the integration of UAVs with LLMs. Academia typically focuses on theoretical modeling and the exploration of new technologies (e.g., 52.7% using External-Offboard Invocation modes), while industry is more concerned with practical applications and system integration (e.g., 50% using Hybrid-Onboard Assistive modes). Researchers should pay more attention to the technological needs in real-world applications, especially regarding real-time performance, robustness, and safety. Researchers should focus on developing lightweight models that can operate with low latency and high efficiency (Li et al., 2025a), particularly for complex tasks such as Multi-UAV Task Reasoning and Planning and Control Logic and Command Generation, optimizing algorithms to improve the accuracy and performance. Additionally, industry developers are concerned with the security and interpretability of large models. Researchers should enhance studies on model transparency and interpretability, developing technologies that ensure data privacy and security in critical tasks, thereby better supporting the widespread application of UAVs.
@@
For practitioners. This study clarifies the tasks that current LLM technologies can support more reliably, such as task planning, user interaction, and automatic report generation. These tasks show stronger performance and lower implementation risks. The comparative analysis also indicates that many complex academic solutions are not yet ready for real-world use because of issues with latency, cost, and reliability. The results further point out important challenges in real-time processing and computational efficiency, giving developers a more realistic view of short-term priorities. To this end, practitioners can move forward in two directions. First, they can apply LLMs in tasks that are already feasible, such as building more natural user interfaces, generating reports automatically, and supporting mission planning. These areas involve lower risks and can deliver practical value relatively quickly. Second, for tasks that are not yet suitable for deployment, developers can focus on researching and testing improvements. This includes integrating lighter models, developing smaller or domain-specific models, or combining LLMs with traditional UAV control systems. In this way, practitioners can benefit from the current capabilities of LLMs while also preparing step by step for more advanced applications in the future.
6.2. Challenges of LLM–UAV Applications
@@
(1) Bridging the Academia-Industry Gap in UAV-LLM Integration: There is a mismatch between academic research and industrial solutions. According to our survey, 67.3% of practitioners believed that academia and industry focus on different directions, and 63.5% thought that their technical requirements differ. The main reason is that academic research often concentrates on theoretical problems and overlooks industrial needs for real-time performance, stability, and system integration (75%). Industry, on the other hand, focuses on commercialization and rapid deployment, which slows down the transfer of academic outcomes to practice. To accelerate technology transfer, future research should better align with industrial needs. Researchers can collaborate with industry to obtain practical data and feedback, ensuring that new technologies can be effectively deployed in real-world environments and avoiding unnecessary delays and obstacles during technology transfer.
@@
(2) Improving safety to increase practitioner trust in LLMs: Deploying LLMs on UAVs faces obstacles related to safety and reliability. Our survey shows that 54.8% of teams believe existing solutions are sufficient, while 45.2% highlight the security risks of large models. One practitioner explicitly stated: “LLMs should not be on a critical system due to their ‘black box’ nature. They are largely unneeded also.” These concerns stem from the lack of interpretability and transparency in LLMs, which can lead to uncertainty in decision-making and increase the risk of system failures. To improve the suitability of LLMs for UAV systems, future research should focus on enhancing interpretability and transparency, developing models that can explain their decision processes. In addition, developers should be provided with effective tools to monitor, validate, and debug model behavior, ensuring alignment with real task requirements, especially in complex and high-risk scenarios.
(3) Enhancing LLM Capabilities for Complex and Robust Autonomous Operations: Current LLMs are not sufficient to handle high-risk and long-duration autonomous UAV missions. As observed in Finding 8, industry interest in using LLMs for direct control execution is low (score: 2.05), with a strong preference for supportive roles (82.7%). Some practitioners further emphasized that “AI should never be put in direct control of a uav.” This reflects a reliance on traditional control algorithms rather than LLMs for decision-making, since current LLMs are viewed as unable to meet performance requirements (51.6%), and their capability in control logic generation remains immature (lowest score: 2.51). Although large-scale LLMs offer advantages in many complex tasks, their high computational resource demands limit their use in UAV systems, especially those requiring real-time performance with hardware constraints. One practitioner noted: “Lightweight large model research still lacks clear quantitative benchmarks for key trade-offs, so careful consideration is needed in model selection and distillation.” Therefore, future work should focus on developing efficient, lightweight models with lower computational demands to meet the real-time processing needs of UAV platforms. In addition, lightweight models should also be designed to handle the computational requirements of complex tasks such as multi-UAV coordination and task planning, ensuring stable and efficient operation in large-scale missions.
6.3. Threats To Validity
6.3.1. Internal Validity
@@
During the process of literature screening and classification, researchers’ subjective judgments may influence the inclusion criteria and labeling of studies, potentially introducing bias into the analysis. To mitigate this threat, we implemented a dual-reviewer protocol across all stages of the analysis. Two authors independently conducted literature selection and labeling, and any disagreements were thoroughly discussed and resolved through consensus. It is worth noting that both authors involved in this process have more than two years of research experience in the field of UAVs, which helps ensure the validity and consistency of the analysis. In the survey component, practitioners may have provided dishonest or unprofessional answers. To reduce this risk, we explicitly informed participants at the beginning of the questionnaire that no personally identifiable information would be collected and that the survey was conducted anonymously. Except for the first four questions, all others were multiple-choice items, with the option to select “I don’t understand this question” for unclear items. Practitioners were free to skip any questions they found ambiguous. This design aimed to improve data quality while protecting practitioners’ privacy. According to Ong and Weiss (2000) (Ong and Weiss, 2000), ensuring anonymity and confidentiality in surveys is effective in eliciting unbiased responses. To collect more responses and overcome language barriers, we translated the questionnaire into Chinese and distributed it via platforms accessible in mainland China. While we made every effort to ensure consistency between the English and Chinese versions, minor discrepancies may have arisen during translation, which could affect practitioners’ understanding. Moreover, as all responses were written in Chinese, translation into English was required during data analysis, potentially introducing interpretation errors. To minimize this risk, two Chinese authors with strong English proficiency collaboratively reviewed and repeatedly refined the translations to ensure their accuracy and consistency.
6.3.2. External Validity
[…]
7. Related Work
7.1. UAV-LLM Integration

### v2

@@ 5.2.1 Practices and Challenges in UAV-LLM integration
Although LLMs show strong potential in auxiliary tasks, their stability, interpretability, and real-time performance have not yet met the high standards required by developers for key tasks such as flight control in real-world UAV projects. Therefore, LLMs are currently more applied in low-risk scenarios such as Task Planning and decision support, while their use in high-risk areas like flight control remains highly restricted. One developer in the survey pointed out, “I do not think LLMs are needed for tasks inboard the uav. AI can have a place on board for image processing to add sensors. But AI should never be put in direct control of a uav. It is an unnecessary, lazy, and non-robust approach to vehicle control.” This clearly reflects the developers’ cautious attitude towards the use of LLMs in critical UAV control tasks, believing that LLMs are better suited to supporting tasks rather than directly participating in flight control.
@@
The Need of UAV-LLM integration. Additionally, we asked developers about the level of need for LLMs in different UAV tasks. The results show that developers have a higher demand for Information Input (score 2.42) and Task Planning (score 2.53). This indicates that LLMs have significant application potential in processing user command translation, perception data, and Task Planning. These tasks require high data processing, reasoning, and planning capabilities, and therefore, there is a strong demand for LLMs’ involvement in these areas. In contrast, the demand for Control Execution (score 2.05) is lower, suggesting that in the areas of flight control and command generation, existing techniques are already well-suited to meet the needs, or the current LLMs cannot be fully relied upon. This indicates that in real-time flight control systems, traditional control algorithms and hardware systems still dominate.
[…]
5.2.2 Researchers-Developers Comparison.
@@
Developer’s Perception. To understand the practical application of academic research in industrial settings, we asked developers whether they refer to academic research in their actual work and included an open-ended question asking for the reasons. The results show that 69.2% of developers indicated that they do refer to academic research. The main reasons for this are as follows: gaining cutting-edge knowledge, verifying the accuracy of information, accelerating development speed, and learning advanced algorithms and approaches. In particular, in the area of algorithm development, academic research provides developers with new ideas and methods, which have actively contributed to improving work efficiency and development quality. For example, one developer mentioned, “I am implementing algorithms for paths’ planning, etc., from the academic publications, I also use them during brainstorming of new ideas and to learn state-of-the-art.
[…]
@@
Although most developers tend to refer to academic research, a small portion of developers do not directly refer to academic outcomes. In this regard, we also asked for their views on the differences between academia and industry in the process of integrating UAVs with large models. According to the survey results, 51.9% of developers believe that the industry is progressing faster than academia. This suggests that while academia is advancing in theoretical research and exploring new approaches, the actual application and implementation in industry are happening at a faster pace. In contrast, 36.5% of developers believe that academia is progressing faster than industry. This reflects that, in basic theoretical research and innovative methods, academia’s breakthroughs provide new ideas and directions for the industry. However, because academic research is not always aligned with industrial demands, the practical application of these progresses relatively slowly. For example, one developer stated, “Industry is faster than academia at some things that lead to higher technology readiness levels. Industry tends not to be good at starting low trl development. In my experience, academia often develops things with little to no industrial application.” This highlights the advantage of industry in applying technology, especially in fields with higher Technology Readiness Levels (TRLs), where industry’s pace of advancement is faster. However, academia often conducts research at low TRL stages, leading to a disconnect between certain academic studies and practical industrial applications.
@@
Additionally, 38.5% of developers believe that there are differences between academia and industry in the integration methods of LLMs and UAVs. 67.3% of developers think that academia and industry focus on different directions. Some developers pointed out: “Research - one topic, but deep, with logical knowledge base and understanding, industry - a lot of topics skimmed, practical usage with ‘black-box-magic”’. These statements illustrate that researchers typically focus on deeply solving specific problems and exploring unknowns and theoretical innovations, while developers are more concerned with practical applications, focusing on how to apply quickly, efficiently, and reliably.
Table 3: Developers’ perceptions on the reasons for differences
@@
Technical Maturity Actual projects require mature and stable technique (75.0%) 84.6%
Difficulty in actual project development and integration (55.8%)
@@
Solutions Desired by Developers. Furthermore, we asked developers to identify the reasons and solutions for these differences. The responses can be categorized into the five challenges listed in  3. Among them, 84.6% of developers identified “Technical Maturity” as the main challenge. This indicates that, unlike academic research, which often emphasizes theoretical innovation, industry places greater importance on mature and reliable technical solutions. Specifically, developers pointed to the need to address issues such as “Actual projects require mature and stable technique” (75.0%), “Difficulty in actual project development and integration” (55.8%), and “Immaturity of lightweight large models” (40.4%). These findings reflect the challenges academic research faces in translating results into real-world applications, as industry often encounters complex conditions during project development and integration.
Developers also noted that industry faces challenges related to “Performance” (48.1%), “Safety” (42.3%), “Cost” (44.2%), and “Other” (57.7%). Corresponding issues include “High computational resource requirements” (46.2%), “Conflicts between latency and real-time requirements” (34.6%), “High computation/hardware costs” (32.7%), ”Model robustness and reliability challenges” (30.8%), and “Commercial return on investment” (28.8%). One developer emphasized: “Industry wants practical results fast, especially smaller startups, etc. Academia has smaller groups to work with, less money, but great research skills and expert knowledge (even if not always practical).” This highlights the urgent demand for rapid results in industry, particularly in the process of commercialization, which makes the transfer of academic research outcomes into practical industrial applications more challenging.
@@ 6 Discussion
This section first explains the significance of this study for researchers and developers, then summarizes the three main challenges of UAV–LLM applications, and finally presents the validity threats.
6.1 Implications
@@
For researchers. This study reveals the significant differences between researchers’ and developers’ practices in integrating UAVs with LLMs. Researchers typically focus on theoretical modeling and the exploration of new approaches (e.g., 52.7% using External-Offboard Invocation modes), while developers are more concerned with practical applications and system integration (e.g., 50% using Hybrid-Onboard Assistive modes). Researchers should pay more attention to the needs in real-world applications, especially regarding real-time performance, robustness, and safety. Researchers should focus on developing lightweight models that can operate with low latency and high efficiency on resource-constrained devices [edgeLLM2025survey, xu2024ondevicelanguagemodelscomprehensive], particularly for complex tasks such as Multi-UAV Task Reasoning and Planning and Control Logic and Command Generation. Additionally, industry developers are concerned with the security and interpretability of large models. Researchers should enhance studies on Explainable AI (XAI) [Black-Box] and trust mechanisms. Developing approaches that ensure data privacy and model transparency in critical tasks, thereby better supporting the widespread application of UAVs.
@@
For developers. This study clarifies the tasks that current LLMs can reliably support, such as task planning, user interaction, and automatic report generation. These tasks show stronger performance and lower implementation risks. The comparative analysis also indicates that many complex academic solutions are not yet ready for real-world use because of issues with latency, cost, and reliability. The results further point out important challenges in real-time processing and computational efficiency, giving developers a more realistic view of short-term priorities. To this end, developers can move forward in two directions. First, they can apply LLMs in tasks that are already feasible, such as building more natural user interfaces, generating reports automatically, and supporting mission planning. These areas involve lower risks and can deliver practical value relatively quickly. Second, for tasks that are not yet suitable for deployment, developers can focus on researching and testing improvements. This includes integrating lighter models [xu2024ondevicelanguagemodelscomprehensive], developing smaller or domain-specific models, or combining LLMs with traditional UAV control systems [brunke2022safe]. In this way, developers can benefit from the current capabilities of LLMs while also preparing step by step for more advanced applications in the future.
6.2 Future Directions for LLM–UAV Applications
@@
(1) Bridging the Academia-Industry Gap in UAV-LLM Integration: There is a mismatch between academic research and industrial solutions. According to our survey, 67.3% of developers believed that academia and industry focus on different directions, and 63.5% thought that their technical requirements differ. The main reason is that researchers often concentrate on theoretical problems and overlook the real-world needs of UAV projects for real-time performance, stability, and system integration (75%). Developers, on the other hand, focus on commercialization and rapid deployment, which slows down the transfer of academic outcomes to practice. To accelerate technology transfer, future research should better align with developers’ needs. Researchers can collaborate with industry to obtain practical data and feedback, ensuring new technologies are effectively deployed in real-world environments and avoiding unnecessary delays and obstacles during technology transfer.
@@
(2) Improving the Safety and Reliability of LLMs to Enhance Developer Trust: Deploying LLMs on UAVs faces obstacles related to safety and reliability. Our survey shows that 45.2% of developers highlight the security risks posed by large models, while 82.7% prefer using LLMs in supportive roles. Their interest in using LLMs for direct control execution is low (score: 2.05). Some developers further emphasized that “LLMs should not be on a critical system due to their ‘black box’ nature” and that “AI should never be put in direct control of a UAV.” On the one hand, these concerns stem from the lack of interpretability and transparency in LLMs, which can lead to uncertainty in decision-making and increase the risk of system failures. On the other hand, they reflect a reliance on traditional control algorithms rather than LLMs for decision-making, since current LLMs are viewed as unable to meet performance requirements (51.6%) and their capability in control logic generation remains immature (lowest score: 2.51). To improve the suitability of LLMs for UAV systems, future research should enhance their interpretability and transparency by developing models that explain their decision-making processes, thereby increasing developers’ trust in LLMs. In addition, it should also develop efficient, lightweight models with lower computational demands to meet the real-time processing and computational requirements of complex tasks on UAV platforms.
6.3 Threats To Validity
6.3.1 Internal Validity
@@
During the process of literature screening and classification, researchers’ subjective judgments may influence the inclusion criteria and labeling of studies, potentially introducing bias into the analysis. To mitigate this threat, we implemented a dual-reviewer protocol across all stages of the analysis. Two authors independently conducted literature selection and labeling, and any disagreements were thoroughly discussed and resolved through consensus. It is worth noting that both authors involved in this process have more than two years of research experience in the field of UAVs, which helps ensure the validity and consistency of the analysis. In the survey component, developers may have provided dishonest or unprofessional answers. To reduce this risk, we explicitly informed participants at the beginning of the questionnaire that no personally identifiable information would be collected and that the survey was conducted anonymously. Except for the first four questions, all others were multiple-choice items, with the option to select “I don’t understand this question” for unclear items. Developers were free to skip any questions they found ambiguous. This design aimed to improve data quality while protecting developers’ privacy. According to Ong and Weiss (2000) [ong2000impact], ensuring anonymity and confidentiality in surveys is effective in eliciting unbiased responses. To collect more responses and overcome language barriers, we translated the questionnaire into Chinese and distributed it via platforms accessible in mainland China. While we made every effort to ensure consistency between the English and Chinese versions, minor discrepancies may have arisen during translation, which could affect developers’ understanding. Moreover, as all responses were written in Chinese, translation into English was required during data analysis, potentially introducing interpretation errors. To minimize this risk, two Chinese authors with strong English proficiency collaboratively reviewed and repeatedly refined the translations to ensure their accuracy and consistency.
6.3.2 External Validity
This study collected valid survey responses by sending invitation emails to active developers on GitHub. However, it is possible that certain groups of developers, such as those working on proprietary systems or those less active in the open-source community, were not adequately covered, and their perspectives may differ from those represented in our sample. Nevertheless, the survey results indicate that developers exhibit substantial diversity in terms of UAV-related experience, professional roles, and levels of open-source project involvement. This diversity enhances our confidence in the representativeness of the sample and supports the generalizability of the findings. Notably, as UAV platforms continue to evolve and LLMs undergo rapid advancement, future systems may integrate new tasks or adopt optimized architectures to address existing limitations in deploying LLMs on UAVs. Consequently, some of the specific findings and insights presented in this study may become outdated over time. This is an inherent limitation in empirical research conducted in fast-developing technological domains. Despite this, the methodology and task framework proposed in this work remain broadly applicable, and our findings offer a timely and valuable reference for developers and researchers seeking to understand and advance UAV-LLM integration.
7 Related Work
7.1 UAV-LLM Integration